OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Show your work. Sn: 2 + 5 + 8 + . . . + ( 3n - 1) = n(1 + 3n)/2 my answer: S1: 2+5+8+(3*1-1)=1(1+3*1)/2 =2 S2: 2+5+8+(3*2-1)=2(1+3*2)/2 =4 S3: 2+5+8+(3*3-1)=3(1+3*3)/2 =5 All the statements are true. my teacher said: n is the sum of n terms given by the formula. Here Sn = n(1+3n)/2. Use the expression for the term to find each successive term and then show that the sum formula works for these n values, i.e. n=1, 2 & 3. @mathmath333

OpenStudy (anonymous):

i already did what she is asking n=1,2,3 then whats left ? |@mathmath333

OpenStudy (mathmath333):

well i asked u before does ur teacher asked for proof by induction or not

OpenStudy (anonymous):

this is anathor Question

OpenStudy (anonymous):

i dont think so she said: n is the sum of n terms given by the formula. Here Sn = n(1+3n)/2. Use the expression for the term to find each successive term and then show that the sum formula works for these n values, i.e. n=1, 2 & 3.

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\[n(1+3n)/2\\for~n=1\\1(1+3\times 1)/2=2\\for~n=2\\2(1+3\times 2)/2=7\\for~n=3\\3(1+3\times 3)/2=15\]

OpenStudy (anonymous):

please explain a bit

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

u have to calculate the sum upto the nth term suppose u have to calculate upto 3 terms so series would be \(2+5+8=15\) but u can also use the formula \(\dfrac{n(1+3n)}{2}\\ =3\dfrac{(1+3\times 3)}{2}\\=15\)

OpenStudy (anonymous):

both of ur answer is same just deifferent way

OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. Show your work. 2 is a factor of n2 - n + 2

OpenStudy (anonymous):

Steps for a proof by induction:- 1 First prove it is true for n=1 2.Assume it is true for n=k 3. Show that it is true for n= (k+1) This may require you to do some algebraic manipulations but once you complete the last step your proof is done. You will find if you remove the parentheses you can simplify into the original expression for k + another term that is divisible by 2. @mathmath333 @hartnn please help

OpenStudy (anonymous):

thats all i could do: S(n) : 2 is a factor of n^2 S(1) : 2 is a factor of 1^2 S(2) : 2 is a factor of 2^2 are S(1) and S(2) true

OpenStudy (anonymous):

@hartnn dear u there please help

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\color{red}{1.)}\\~\\ for~n=1\\~\\ n^2 - n + 2\\~\\ =(1)^2 - 1 + 2\\~\\ =2\\~\\ \text{hence it is true for }~n=1\\~\\ \color{red}{2.)}\\~\\ \text{assume that it is true for }~n=k\\~\\ \text{so it is true for }\\~\\ k^2 - k + 2\\~\\ \text{now we have to prove it is true for }~n=k+1\\~\\ (k+1)^2 -( k+1) + 2\\~\\ =k^2+2k+1 - k-1 + 2\\~\\ =(k^2 - k+ 2)+2k\\~\\ (k^2 - k+ 2) ~\text{ is true as we assumed }\\~\\ \text{and 2k is a factor of 2 }\\~\\ \text{PROVED! }\\~\\ }\end{align}\)

OpenStudy (anonymous):

thanks =)